# Distributed Discrete-time Dynamic Outer Approximation of the Intersection of Ellipsoids

### v1

@@ Abstract
This paper presents the first discrete-time distributed algorithm to track the tightest ellipsoids that outer approximates the global dynamic intersection of ellipsoids. The ellipsoids are defined as time-varying positive definite matrices. On the other hand, given an undirected network, each node is equipped with one of these ellipsoids. The solution is based on a novel distributed reformulation of the original centralized semi-definite outer Löwner-John program, characterized by a non-separable objective function and global constraints. We prove finite-time convergence to the global minima of the centralized problem in the static case and finite-time bounded tracking error in the dynamic case. Moreover, we prove boundedness of estimation in the tracking of the global optimum and robustness in the estimation against time-varying inputs. As a by-product, the proposed algorithm extends min/max dynamic consensus algorithms to positive definite matrices. We illustrate the properties of the algorithm with different simulated examples, including a distributed estimation showcase where our proposal is integrated into a distributed Kalman filter to surpass the state-of-the-art in mean square error performance.
@@ I Introduction
From safe control to distributed sensor fusion, one fundamental problem in control systems is how to approximate system measurements, states, or constraints such that their essential features are preserved, being simple enough to be handled. One of the most popular representations is ellipsoids [1], characterized by symmetric and positive semi-definite $n$-dimensional matrices. Ellipsoids are chosen to describe the conservative shape of an obstacle [2, 3], obstacle-free regions [4, 5], estimated quantities with their associated uncertainty [6, 7], very large multi-agent populations [8], or patterns or geometrical objects to be identified and clustered [9]. All these examples are related to or can be posed as finding the tightest outer ellipsoid approximating a convex set [10].
@@
In this work, we propose the first distributed discrete-time dynamic algorithm to compute the solution of the original centralized outer Löwner-John method. From a distributed optimization point of view, compared to other works (Sec. II), the Löwner-John problem is comprised by a centralized optimization program with a non-separable objective function over positive definite matrices and with global coupling constraints (Sec. III), preventing the use of current distributed optimization methods. The proposed algorithm is based on a novel distributed reformulation of the outer Löwner-John method (Sec. IV). Each node solves a local semi-definite program whilst the agreement value tracks the dynamic intersection of ellipsoids. We analyze the theoretical properties of the algorithm (Sec. V). For constant input ellipsoids, we prove finite-time convergence to the global minima of the centralized problem; for time-varying input ellipsoids, we prove finite-time bounded tracking error. We prove that our algorithm guarantees boundedness of the estimates and robustness on the feasibility of global tracking under dynamic input ellipsoids. We illustrate the proposed algorithm in simulated experiments (Sec. VII), including an integration on a distributed Kalman filter for stochastic estimation (Sec. VI).
@@ II Related Work
Many distributed optimization solutions depart from a centralized optimization formulation and then derive consensus-based algorithms to reconstruct global quantities and converge to the global minimizer [14]. When the objective function is separable as a sum of local (strongly) convex smooth functions, dynamic consensus over the optimization variables and the gradients [15] can be exploited to design distributed versions of gradient descent [16], stochastic gradient descent [17], general first order optimization methods [18] or second order newton-raphson-like methods [19]. Despite proving the success of consensus-based approaches for distributed optimization, these solutions are not suitable when the objective function cannot be decomposed into local smooth convex functions.
To deal with constraints, some works propose extensions of consensus-based algorithms that rely on projection methods [20], population dynamics equations [21], subgradient methods [22], or recurrent neural networks [23]. The most popular alternative to address global coupling constraints is the use of a primal and/or dual proximal stage at each node [24]. In particular, the alternating direction method of multipliers [25] and the Douglas-Rachford splitting [26] allow to handle equality coupling constraints among nodes by solving simultaneously the primal and dual optimization problem. Typically, these distributed proximal methods either assume separability of the original objective function [27] and/or are restricted to scalar quantities [28]. Therefore, they are not suitable for distributing the outer Löwner-John method. In the presence of constraints and a linear objective function, it is possible to find distributed semi-definite problem reformulations [29] that are solved through primal-dual methods. With a similar spirit to our work, [30] proposes a polyhedral outer approximation in a distributed optimization context. Nevertheless, in this case the polyhedral approximation represents a convex constraint set rather than the optimization objective.
@@ III Problem statement
At each discrete-time step $k\in\{0,1,\dots\}$, each agent $i\in\mathcal{I}$ is endowed with a positive definitive input matrix $\mathbf{P}_{i}[k]\in\mathbb{S}^{n}_{++}$. To provide some intuition, this matrix can be pictured as representing the uncertainty over an estimate [35, 36], where $\mathbf{P}_{i}[k]^{-1}$ represents its information matrix. Such information matrix characterizes the ellipsoid $\mathcal{E}(\mathbf{P}_{i}[k]^{-1})=\{\mathbf{y}\in\mathbb{R}^{n}:\mathbf{y}^{\top}\mathbf{P}_{i}[k]^{-1}\mathbf{y}\leq 1\}$. In practice, the input matrices $\mathbf{P}_{i}[k]$ cannot have arbitrarily large values, and the variation of $\mathbf{P}_{i}[k]$ from instant $k$ to $k+1$ cannot be arbitrarily large, so in this work we do the following assumption.
@@
There exist $\underline{p}\leq\overline{p}$, $0\leq\underline{\theta}\leq 1\leq\bar{\theta}$ such that the input matrices $\mathbf{P}_{i}[k]$, $\forall i\in\mathcal{I}$ and $\forall k\in\{0,1,\dots\}$, satisfy
$\underline{p}\mathbf{I}$ $\preceq\mathbf{P}_{i}[k]\preceq\overline{p}\mathbf{I}$
$\underline{\theta}\mathbf{P}_{i}[k-1]$ $\preceq\mathbf{P}_{i}[k]\preceq\bar{\theta}\mathbf{P}_{i}[k-1]$
The previous inequalities provide bounds for the eigenvalues of $\mathbf{P}_{i}[k]$, absolute for all time and relative to the previous instant respectively.
@@
Figure 1: Problem setting: (top) each node in graph $\mathcal{G}$ is equipped with a time-varying ellipsoid described by the matrix $\mathbf{P}_{i}[k]\quad\forall i\in\{1,2,3,4\}$. The goal is to cooperate to find the smallest ellipsoid that outer approximates their intersection (black ellipsoid for $\mathbf{Q}^{*}[k]$); (bottom) each node (e.g., node $2$) only has access to its neighboring information, leading to conservative approximations of $\mathbf{Q}^{*}[k]$ (black ellipsoid for $\mathbf{Q}_{2}[k]$). In this work we propose a distributed method to make $\mathbf{Q}_{i}[k]$ converge to $\mathbf{Q}^{*}[k]$ using local information.
@@
    $f$ is convex, meaning that, for any $\lambda_{1},\lambda_{2}\geq 0,$ $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\mathbb{S}_{+}^{n}$ with $\lambda_{1}+\lambda_{2}=1$ it follows that: $f(\lambda_{1}\mathbf{Q}_{1}+\lambda_{2}\mathbf{Q}_{2})\leq\lambda_{1}f(\mathbf{Q}_{1})+\lambda_{2}f(\mathbf{Q}_{2})$.
@@
Assumption 2 holds for many popular choices of $f(\bullet)$ related to the volume of ellipsoids or information theoretic measures, such as $-\text{tr}(\bullet),$ $\log(\text{ {det}}((\bullet)^{-1})),$ or $\text{tr}((\bullet)^{-1})$. Moreover, Assumption 2-4) is an instrumental property to be used in subsequent analysis.
The program in (1) is known as the outer Löwner-John method. The formulation of the outer Löwner-John method is centralized, in the sense that all the inputs $\mathbf{P}_{i}[k]$ are needed to compute $\mathbf{Q}^{*}[k]$. However, in the aforementioned cooperative tasks, it is of key importance to develop a distributed solution such that each agent has a local version $\mathbf{Q}_{i}[k]$ of $\mathbf{Q}^{*}[k]$ computed from local interactions.
@@ IV Distributed Discrete-time Outer Ellipse Computation
with initial conditions $\mathbf{Q}_{i}[0]=\mathbf{P}_{i}[0]^{-1}$. Note that each node only requires the solution of (2) at each neighboring node in the previous instant. Therefore, (2) defines a distributed algorithm, as detailed in Algorithm 1.
Algorithm 1 Distributed estimation of the outer Löwner-John ellipsoid at node $i$
@@
The properties of our proposed approach are summarized in the following Theorem.
@@
Let Assumptions 1 and 2 hold and $\mathcal{G}$ be connected. Then, the following statements are true:
@@
    (Robustness) $\mathcal{C}_{i}[k]\subseteq\mathcal{C}^{*}[k]$ for all $k\in\{0,1\dots\},i\in\mathcal{I}$.
@@
    (Boundedness) $\mathbf{Q}_{i}[k]\preceq\underline{p}\mathbf{I}$ for all $k\in\{0,1\dots\},i\in\mathcal{I}$.
@@
    $\|\mathbf{Q}_{i}[k]-\mathbf{Q}^{*}[k]\|\leq\delta$
    for any $k\geq K,i\in\mathcal{I}$, and with $\mathbf{Q}_{i}[k]=\mathbf{Q}^{*}[k]$ if $\underline{\theta}=\bar{\theta}=1$.
@@
First, the intuition behind robustness is that, given arbitrary input matrices $\{\mathbf{P}_{i}[k]\}_{i=1}^{\mathsf{N}}$ under Assumption 1, the feasible set of the local optimization problem at each node is always contained inside the feasible set of the global optimization problem; this is important because, if robustness did not hold, then the optimum of the local optimization problem at node $i$ might be such that $\mathbf{Q}_{i}[k]\notin\mathcal{C}^{*}[k]$ and, therefore, it would not be possible to converge arbitrarily close to the global optimum. For instance, Fig. 2 visualizes a case where, if $f(\mathbf{Q}_{i}[k-1])<f(\mathbf{Q}^{*}[k])$ $\forall k>K$ and robustness did not hold, then $\mathbf{Q}_{i}[k]=\mathbf{Q}_{i}[k-1]$ $\forall k>K$, which means that the estimate at node $i$ would get stuck in a potentially unfeasible point forever. Thus, the proposed algorithm is robust against changes in the input ellipsoids, being able to converge to the global optimum always. Second, boundedness implies that, irrespective of the changes in the input ellipsoids, the estimate of the global optimum at each node never escapes to infinity. Third, convergence means that our algorithm converges in finite time $K$ to a region around the optimum of the global problem, where the size of the region depends on how fast the input ellipsoids vary with time: in particular, if the input ellipsoids are constant, then $\delta=0$ and the global optimum is perfectly recovered; besides, the estimates remain inside that region for $k>K$.
@@ V-A Auxiliary results
[…]
@@
For the first item, since $f(\bullet)$ and $\mathcal{C}_{1},\mathcal{C}_{2}$ are convex, then $\mathbf{Q}_{1},\mathbf{Q}_{2}$ must be unique. Hence, if $\mathbf{Q}_{2}\in\mathcal{C}_{1}$, then $\mathbf{Q}_{1}=\mathbf{Q}_{2}$ must follow. For the second item, $\mathcal{C}_{1}$ is closed, so the relative boundary ${\textsf{rebdr}}(\mathcal{C}_{1})$ exists and contains all its relative boundary points. The next steps of the proof follow by contradiction. Assume that $\mathbf{Q}_{1}\in{\textsf{relint}}(\mathcal{C}_{1})$. Then, there exist a ball $\mathcal{B}\in{\textsf{relint}}(\mathcal{C}_{1})$ centered at $\mathbf{Q}_{1}$ such that $f(\mathbf{Q}_{1})\leq f(\mathbf{Q}),\forall\mathbf{Q}\in\mathcal{B}$. Thus, $\mathbf{Q}_{1}$ is a local optimum of $f(\bullet)$, and, as a result of convexity of $f(\bullet)$ and $\mathcal{C}_{1},\mathcal{C}_{2}$, it is the global optimum of $f(\bullet)$ in any case. By assumption of the lemma, $\mathbf{Q}_{1}\in\mathcal{C}_{1}\subseteq\mathcal{C}_{2}$ and thus $\mathbf{Q}_{1}=\mathbf{Q}_{2}$. However, by assumption of the second item, $\mathbf{Q}_{2}\notin\mathcal{C}_{1}$, which leads to the contradiction $\mathbf{Q}_{1}\notin\mathcal{C}_{1}$. ∎
The next technical lemma demonstrate properties of the feasible set $\mathcal{C}_{i}[k]$ and $\mathcal{C}^{*}[k]$ that are important to study the relationships between the centralized and distributed optimization problems (1) and (2).
@@
    The sets $\mathcal{C}_{i}[k],\mathcal{C}^{*}[k]$ are closed and convex for all $k\in\{0,1,\dots\}$ and $i\in\mathcal{I}$.
@@
For item 3) we proceed by induction. As induction base, set an arbitrary agent $i\in\mathcal{I}$ which holds initial condition $\mathbf{Q}_{i}[0]=\mathbf{P}_{i}^{-1}[0]\in\mathcal{C}_{i}[0]:=\{\mathbf{P}_{i}[k]^{-1}\}$. Now, referring to the centralized optimization problem (1), let $\lambda_{j}=1$ if $j=i$ and $\lambda_{j}=0$ otherwise. Hence, $\mathbf{Q}_{i}[0]=\sum_{j=1}^{\textsf{N}}\lambda_{j}\mathbf{P}_{j}[k]^{-1}\in\mathcal{C}^{*}[k]$. Henceforth, $\mathcal{C}_{i}[0]\subseteq\mathcal{C}^{*}[0]$.
@@
For an arbitrary $\mathbf{Q}\in\mathcal{C}_{i}[k]$ it follows that

### v2

@@ Abstract
This paper presents the first discrete-time distributed algorithm to track the tightest ellipsoids that outer approximates the global dynamic intersection of ellipsoids. Given an undirected network, we consider a setup where each node measures an ellipsoid, defined as a time-varying positive semidefinite matrix. The goal is to devise a distributed algorithm to track the tightest outer approximation of the intersection of all the ellipsoids. The solution is based on a novel distributed reformulation of the original centralized semi-definite outer Löwner-John program, characterized by a non-separable objective function and global constraints. We prove finite-time convergence to the global minima of the centralized problem in the static case and finite-time bounded tracking error in the dynamic case. Moreover, we prove boundedness of estimation in the tracking of the global optimum and robustness in the estimation against time-varying inputs. We illustrate the properties of the algorithm with different simulated examples, including a distributed estimation showcase where our proposal is integrated into a distributed Kalman filter to surpass the state-of-the-art in mean square error performance.
@@
©This paper has been accepted for publication at IEEE Transactions on Automatic Control. Please, when citing the paper, refer to the official manuscript with the following DOI: 10.1109/TAC.2025.3540423.
@@ I Introduction
From safe control to distributed sensor fusion, one fundamental problem in control systems is how to approximate system measurements, states, or constraints such that their essential features are preserved, being simple enough to be handled. One of the most popular representations is ellipsoids [1], characterized by symmetric and positive semi-definite $n$-dimensional matrices. Ellipsoids are chosen to describe the conservative shape of an obstacle [2], obstacle-free regions [3], estimated quantities with their associated uncertainty [4], very large multi-agent populations [5], or patterns and geometrical objects to be identified and clustered [6]. All these examples are related to or can be posed as finding the tightest outer ellipsoid approximating a convex set [7].
@@
In this work, we propose the first distributed discrete-time dynamic algorithm to compute the solution of the original centralized outer Löwner-John method, approximating the global solution to the ellipsoid intersection problem. From a distributed optimization point of view, compared to other works (Sec. II), the Löwner-John problem is comprised by a centralized optimization program with a non-separable objective function over positive definite matrices and with global coupling constraints (Sec. III), preventing the use of current distributed optimization methods. The proposed algorithm is based on a novel distributed reformulation of the outer Löwner-John method (Sec. IV). Each node solves a local semi-definite program whilst the agreement value tracks the dynamic intersection of ellipsoids. We analyze the theoretical properties of the algorithm (Sec. V). For constant input ellipsoids, we prove finite-time convergence to the global minima of the centralized problem; for time-varying input ellipsoids, we prove finite-time bounded tracking error. We prove that our algorithm guarantees boundedness of the estimates and robustness on the feasibility of global tracking under dynamic input ellipsoids. We illustrate the proposed algorithm in simulated experiments (Sec. VII), including an integration on a distributed Kalman filter for stochastic estimation (Sec. VI).
@@ II Related Work
Many distributed optimization solutions depart from a centralized optimization formulation and then derive consensus-based algorithms to reconstruct global quantities and converge to the global minimizer [12]. When the objective function is separable as a sum of local (strongly) convex smooth functions, dynamic consensus over the optimization variables and the gradients [13] can be exploited to design distributed versions of gradient descent [14], general first order optimization methods [15] or second order Newton-Raphson-like methods [16]. Despite proving the success of consensus-based approaches for distributed optimization, these solutions are not suitable when the objective function cannot be decomposed into local smooth convex functions.
To deal with constraints, some works propose extensions of consensus-based algorithms that rely on projection methods [17], population dynamics equations [18], or subgradient methods [19]. The most popular alternative to address global coupling constraints is the use of a primal and/or dual proximal stage at each node [20]. In particular, the alternating direction method of multipliers [21] and the Douglas-Rachford splitting [22] allow to handle equality coupling constraints among nodes by solving simultaneously the primal and dual optimization problem. Typically, these distributed proximal methods either assume separability of the original objective function [23] and/or are restricted to scalar quantities [24]. Therefore, they are not suitable for distributing the outer Löwner-John method. In the presence of constraints and a linear objective function, it is possible to find distributed semi-definite problem reformulations [25] that are solved through primal-dual methods. With a similar spirit to our work, [26] proposes a polyhedral outer approximation in a distributed optimization context. Nevertheless, in this case the polyhedral approximation represents a convex constraint set rather than the optimization objective.
@@
From an application point of view, the outer Löwner-John method has been widely used for stochastic distributed estimation under the name of Covariance Intersection (CI) [31]. CI is a method to consistently integrate estimates of neighboring nodes in networks with unknown cross-correlations. CI is optimal for two nodes [32], but in general it is suboptimal [33]. The consistency property motivates the use of CI for distributed Kalman filtering [34, 35]. Nevertheless, despite improved performance upon classical distributed Kalman filters [36], the lack of optimality guarantees in the fusion leads to suboptimality for these filtering alternatives. Other extensions of CI include the Covariance Union (CU) [37] and the Split Covariance Intersection (SCI) [38]; the former determines the tightest ellipsoid that guarantees consistency regardless of which of the estimates is consistent, while the latter splits the admissible set of cross-covariances to tighten the conservative bound provided by CI. They have both being applied in filtering problems [39, 40]. In pursuit of optimality guarantees in stochastic distributed estimation, we recently used the original outer Löwner-John method to derive the certifiable optimal distributed Kalman filter under unknown correlations [9, 4]. These works restrict the outer Löwner-John method to the local neighborhood of each node, rather than reconstructing the global optimum provided by all the node estimates. For this reason, and to demonstrate the benefits of our proposed algorithm, we integrate the solution in a distributed Kalman filter.
@@ III Problem statement
Consider a network of N agents which communicate according to an undirected graph $\mathcal{G}=(\mathcal{I},{\mathcal{F}})$, where =$\mathcal{I}=\{1,\dots,{\textsf{N}}\}$ is the vertex set and ${\mathcal{F}}\subset\mathcal{I}\times\mathcal{I}$ is the edge set representing the communication links between agents. The set of neighbors of agent $i$, including $i$, is $\mathcal{N}_{i}=\{j\in\mathcal{I}|(i,j)\in{\mathcal{F}}\}\cup\{i\}$.
At each discrete-time step $k\in\{0,1,\dots\}$, each agent $i\in\mathcal{I}$ measures a positive definitive input matrix $\mathbf{P}_{i}[k]\in\mathbb{S}^{n}_{++}$. To provide some intuition, this matrix can be pictured as representing the uncertainty over an estimate [36, 41], where $\mathbf{P}_{i}[k]^{-1}$ represents its information matrix. Such information matrix characterizes the ellipsoid $\mathcal{E}(\mathbf{P}_{i}[k]^{-1})=\{\mathbf{y}\in\mathbb{R}^{n}:\mathbf{y}^{\top}\mathbf{P}_{i}[k]^{-1}\mathbf{y}\leq 1\}$. In practice, the input matrices $\mathbf{P}_{i}[k]$ cannot have arbitrarily large values, and the variation of $\mathbf{P}_{i}[k]$ from instant $k$ to $k+1$ cannot be arbitrarily large, so in this work we do the following assumption.
@@
There exist $0<\underline{p}\leq\overline{p}$ such that the input matrices $\mathbf{P}_{i}[k]$, $\forall i\in\mathcal{I}$ and $\forall k\in\{0,1,\dots\}$, satisfy
$\underline{p}\mathbf{I}$ $\preceq\mathbf{P}_{i}[k]\preceq\overline{p}\mathbf{I}.$
The previous inequalities provide bounds for the eigenvalues of $\mathbf{P}_{i}[k]$ for all time.
@@
Figure 1: Problem setting: (top) each node in graph $\mathcal{G}$ represents a time-varying ellipsoid described by the matrix $\mathbf{P}_{i}[k]\quad\forall i\in\{1,2,3,4\}$. The goal is to cooperate to find the smallest ellipsoid that outer approximates their intersection (black ellipsoid for $\mathbf{Q}^{*}[k]$); (bottom) each node (e.g., node $2$) only has access to its neighboring information, leading to conservative approximations of $\mathbf{Q}^{*}[k]$ (black ellipsoid for $\mathbf{Q}_{2}[k]$). In this work we propose a distributed method to make $\mathbf{Q}_{i}[k]$ converge to $\mathbf{Q}^{*}[k]$ using local information.
@@
    $f$ is strictly convex, meaning that, for any $c_{1},c_{2}{\geq}0,$ $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\mathbb{S}_{+}^{n}$ with $c_{1}+c_{2}=1$ it follows that: $f(c_{1}\mathbf{Q}_{1}+c_{2}\mathbf{Q}_{2})<c_{1}f(\mathbf{Q}_{1})+c_{2}f(\mathbf{Q}_{2})$.
@@
Assumption 2 holds for many popular choices of $f(\bullet)$ related to the volume of ellipsoids or information theoretic measures, such as $\log(\text{ {det}}((\bullet)^{-1}))$ or $\text{tr}((\bullet)^{-1})$. Moreover, Assumption 2-4) is an instrumental property to be used in subsequent analysis.
The program in (1) is known as the outer Löwner-John method and the intuition behind its structure is the following. The minimization of the objective function seeks to minimize the size of the output ellipsoid. For a problem with $\mathsf{N}=2$ nodes, the first constraint of $\mathcal{C}^{*}[k]$ means that, if $\lambda_{1}=0$, then $\mathcal{E}(\mathbf{Q}_{1})\subseteq\mathcal{E}(\mathbf{Q}_{2})$, i.e., the ellipsoid at node $1$ is contained in the ellipsoid at node $2$. In general, $\lambda_{1},\lambda_{2}>0$, which means that the two ellipsoids have an overlapping area in common, characterized by the convex combination of $\mathbf{P}_{1}[k]^{-1},\mathbf{P}_{2}[k]^{-1}$, which is the second constraint in $\mathcal{C}^{*}[k]$. The ellipsoid $\mathbf{Q}^{*}[k]$ is the tightest ellipsoid that covers such overlapping area. The same reasoning extends to the $\mathsf{N}>2$ case.
The formulation of the outer Löwner-John method is centralized, in the sense that all the inputs $\mathbf{P}_{i}[k]$ are needed to compute $\mathbf{Q}^{*}[k]$. However, in the aforementioned cooperative tasks, it is of key importance to develop a distributed solution such that each agent has a local version $\mathbf{Q}_{i}[k]$ of $\mathbf{Q}^{*}[k]$ computed from local interactions.
@@ IV Distributed Discrete-time Outer Ellipse Computation
for $k\geq 1$, with initial conditions $\mathbf{Q}_{i}[0]=\mathbf{P}_{i}[0]^{-1}$ and a parameter $\overline{\theta}$ to be specified subsequently. Note that each node only requires the solution of (2) at each neighbor in the previous instant. Therefore, (2) defines a distributed algorithm, described in Algorithm 1. Its properties are summarized in Theorem 1.
Algorithm 1 Distributed estimation of (1) at node $i$
@@
Let Assumptions 1 and 2 hold and $\mathcal{G}$ be connected. Then, for a network that executes Algorithm 1, the following statements are true:
@@
    (Robustness) $\mathcal{C}_{i}[k]\subseteq\mathcal{C}^{*}[k]$ for all $k\in\{{1,2}\dots\},i\in\mathcal{I}$.
@@
    (Boundedness) $\mathbf{Q}_{i}[k]\preceq{(1/\underline{p})}\mathbf{I}$ for all $k\in\{{1,2}\dots\},i\in\mathcal{I}$.
@@
    ${\|\mathbf{Q}_{i}[k]-\mathbf{Q}^{*}[k]\|\leq\delta\text{ for any }k\geq K,i\in\mathcal{I}.}$
    In addition, if $\underline{\theta}=\bar{\theta}=1$, then $\mathbf{Q}_{i}[k]=\mathbf{Q}^{*}[k]$.
@@
We prove Theorem 1 in Section V-C, after introducing auxiliary results in Sections V-A and V-B. Before that, we describe the main properties expressed in the theorem. The intuition behind robustness is that, given arbitrary input matrices $\{\mathbf{P}_{i}[k]\}_{i=1}^{\mathsf{N}}$ under Assumption 1, the feasible set of the local optimization problem at each node is always contained inside the feasible set of the global optimization problem; this is important because, if robustness did not hold, then the optimum of the local optimization problem at node $i$ might be such that $\mathbf{Q}_{i}[k]\notin\mathcal{C}^{*}[k]$ and, therefore, it would not be possible to converge arbitrarily close to the global optimum. For instance, Fig. 2 visualizes a case where, if $f(\mathbf{Q}_{i}[k-1])<f(\mathbf{Q}^{*}[k])$ $\forall k>K$ and robustness did not hold, then $\mathbf{Q}_{i}[k]=\mathbf{Q}_{i}[k-1]$ $\forall k>K$, which means that the estimate at node $i$ would get stuck in a potentially unfeasible point forever. Thus, the proposed algorithm is robust against changes in the input ellipsoids, being able to converge to the global optimum always. Second, boundedness implies that, irrespective of the changes in the input ellipsoids, the estimate of the global optimum at each node never escapes to infinity. Third, convergence means that our algorithm converges in finite time $K$ to a region around the optimum of the global problem, where the size of the region depends on how fast the input ellipsoids vary with time: in particular, if the input ellipsoids are constant, then $\delta=0$ and the global optimum is perfectly recovered; besides, the estimates remain inside that region for $k>K$.
@@ V-A Auxiliary results
$\mathbf{Q}_{1}=\operatorname*{arg\,min}_{\mathbf{Q}\in\mathcal{C}_{1}}f(\mathbf{Q})\qquad\text{and}\qquad\mathbf{Q}_{2}=\operatorname*{arg\,min}_{\mathbf{Q}\in\mathcal{C}_{2}}f(\mathbf{Q}).$
@@
For the first item, since $f(\bullet)$ is strictly convex and $\mathcal{C}_{1},\mathcal{C}_{2}$ are convex, then $\mathbf{Q}_{1},\mathbf{Q}_{2}$ must be unique. Hence, if $\mathbf{Q}_{2}\in\mathcal{C}_{1}$, then $\mathbf{Q}_{1}=\mathbf{Q}_{2}$ must follow. For the second item, $\mathcal{C}_{1}$ is closed, so the relative boundary ${\textsf{rebdr}}(\mathcal{C}_{1})$ exists and contains all its relative boundary points. The next steps of the proof follow by contradiction. Assume that $\mathbf{Q}_{1}\in{\textsf{relint}}(\mathcal{C}_{1})$. Then, there exist a ball $\mathcal{B}\in{\textsf{relint}}(\mathcal{C}_{1})$ centered at $\mathbf{Q}_{1}$ such that $f(\mathbf{Q}_{1})\leq f(\mathbf{Q}),\forall\mathbf{Q}\in\mathcal{B}$. Thus, $\mathbf{Q}_{1}$ is a local optimum of $f(\bullet)$, and, as a result of convexity of $f(\bullet)$ and $\mathcal{C}_{1},\mathcal{C}_{2}$, it is the global optimum of $f(\bullet)$ in any case. By assumption of the lemma, $\mathbf{Q}_{1}\in\mathcal{C}_{1}\subseteq\mathcal{C}_{2}$ and thus $\mathbf{Q}_{1}=\mathbf{Q}_{2}$. However, by assumption of the second item, $\mathbf{Q}_{2}\notin\mathcal{C}_{1}$, which leads to the contradiction $\mathbf{Q}_{1}\notin\mathcal{C}_{1}$. ∎
The next lemma demonstrates properties of the feasible set $\mathcal{C}_{i}[k]$ and $\mathcal{C}^{*}[k]$ that are important to study the relationships between the centralized and distributed optimization problems (1) and (2).
@@
    The sets $\mathcal{C}_{i}[k],\mathcal{C}^{*}[k]$ are closed and convex for all $k\in\{{1,2},\dots\}$ and $i\in\mathcal{I}$.
@@
For item 3) we proceed by induction. As induction base, set an arbitrary agent $i\in\mathcal{I}$ and define $\mathcal{C}_{i}[0]$ only for $k=0$ as $\mathcal{C}_{i}[0]:=\{\mathbf{P}_{i}[{0}]^{-1}\}$, so that the initial condition comply $\mathbf{Q}_{i}[0]=\mathbf{P}_{i}^{-1}[0]\in\mathcal{C}_{i}[0]$. Now, referring to the centralized optimization problem (1), let $\lambda_{j}=1$ if $j=i$ and $\lambda_{j}=0$ otherwise. Hence, $\mathbf{Q}_{i}[0]=\sum_{j=1}^{\textsf{N}}\lambda_{j}\mathbf{P}_{j}[{0}]^{-1}\in\mathcal{C}^{*}[{0}]$. Henceforth, $\mathcal{C}_{i}[0]\subseteq\mathcal{C}^{*}[0]$.
@@
For an arbitrary $\mathbf{Q}\in\mathcal{C}_{i}[k],{k=1,2,\dots}$ it follows that